OpenStudy (anonymous):

A quadratic function and an exponential function are graphed below. Which graph most likely represents the quadratic function?

OpenStudy (anonymous):

Graph and Answers:

OpenStudy (anonymous):

I think its the second one but im not sure

OpenStudy (anonymous):

the graph is so small it's so hard to differentiate the lines

OpenStudy (anonymous):

@amistre64 can you help?

OpenStudy (anonymous):

@iGreen @aaronq ?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@lilshane

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Exponential function always grows faster than quadratic function, so C makes sense.

OpenStudy (anonymous):

Thank you, @geerky42 , do you mind if I as you two more questions?

geerky42 (geerky42):

Sure

OpenStudy (anonymous):

The table below shows the values of f(x) and g(x) for different values of x. One of the functions is a quadratic function, and the other is an exponential function. Which function is most likely increasing quadratically?

OpenStudy (anonymous):

f(x), because it grows more slowly than g(x) f(x), because it eventually intersects g(x) g(x), because it grows faster than f(x) g(x), because it will not intersect f(x)

geerky42 (geerky42):

Again, exponential function always grows faster than quadratic function. So can you tell which grow slower?

OpenStudy (anonymous):

I feel like f(x) is exponential because it isn't consistent

OpenStudy (anonymous):

and it's growing way faster

geerky42 (geerky42):

Hmm, g(x) seem to grow by being multiplied by 2. 2*2=4; 4*2 = 8, 8*2 = 16, and so on

geerky42 (geerky42):

\(g(x)\) behaves like \(2^x\)

geerky42 (geerky42):

Right?

OpenStudy (anonymous):

I believe so

geerky42 (geerky42):

So we can guess that g is exponential function. So we are left with A or B.

geerky42 (geerky42):

intersections don't really tell us much, so we better off going with A.

OpenStudy (anonymous):

Thank you!